OpenStudy (lime):

What is the distance between P(–4, 3) and Q(6, 1)?

Parth (parthkohli):

Use the distance formula., if you know it!

Parth (parthkohli):

\[distance = {\sqrt(x_2 - x_1)^2 + (y_2 - y_1)^2}\]

OpenStudy (lime):

\[d = \sqrt(x^2 - x^1)^2 + (y^2 - y^1)^2\]

Parth (parthkohli):

Yeah, but 2 and 1 are subs... not exponents!

OpenStudy (lime):

\[d = \sqrt(3 - 1)^2 + (-4 + 6)^2?\]

OpenStudy (lime):

\[= \sqrt(2)^2 + (10)^2\]

OpenStudy (lime):

\[\sqrt(4 + 100) = \sqrt(104) \approx \]

Parth (parthkohli):

Yes

OpenStudy (lime):

And rounded to the nearest tenth is approximately 10.2? Thanks @ParthKohli ! :)

Parth (parthkohli):

Actually you can express it in radical form, since that's more accurate!

Parth (parthkohli):

104|2 52|2 26|2 13|13 -------

Parth (parthkohli):

\[\sqrt{104 } \implies \sqrt{2 \times 2 \times 2 \times 13} \implies 2\sqrt{26}\]